OpenStudy (anonymous):

is 9m^2-1 factorable

OpenStudy (anonymous):

i dont think so

OpenStudy (anonymous):

well actually you could by 1 or -1 but its not necessary to solve/simplify

OpenStudy (anonymous):

(m+1/3)(m-1/3)

OpenStudy (anonymous):

\[9m^2-1 = m^2-1/9\]

OpenStudy (anonymous):

like heck it does.

OpenStudy (amistre64):

yes it is; its a diff of squares

OpenStudy (anonymous):

the right hand side is one ninth of the left. \[(3m-1)(3m+1)=9m^2-1\]

OpenStudy (amistre64):

(3m-1)(3m+1) ; very sticky ... lol

OpenStudy (anonymous):

oh i totally knew that